Can you please answer all of them?

Section B: Free Response (you must show your work for full credit) 1. Draw Frost circle diagrams for each compound shown. Ass4. Provide a plausible mechanism for the reaction shown.(6 pts) de 5. Each transformation shown below requires at least two sProvide a synthesis for the target molecule using benzene. OH SM TM
Answer #1

1.

Anti-bonding ... Non-bonding Bonding The given molecule contains only 1 pi bond and mo negative charges. Hence number of elec

2.

Huckel's number of pi electrons means the electrons which satisfy huckel's rule. This include all the electrons that are in delocalisation in the given molecule.

In our system, 6 pi bonds and 1 lone pair of electron from oxygen are in delocalisation. Since, 6 pi bonds contain 12 (= 6 × 2) electrons and 1 lone pair contains 2 electrons, total number of huckel's pi electrons are 12 + 2 = 14.

Since, it has 6 pi bonds, it will have 6 pi electron pairs.

Also, since huckel's electrons are 14, it follows huckel's rule for aromaticity (4n + 2 pi electrons, where n =3).

3.

In friedel-crafts alkylation, an electrophile attacks the benzene ring. For better reactivity the ring must be electron rich. This means that the ring should contain electron releasing groups.

Electron releasing groups increases the reactivity of friedel-crafts alkylation.

Electron withdrawing groups decreases the reactivity of friedel-crafts alkylation.

In the given compounds,

Br. ENG Br EWG To (B) ENG EDG via I I effect Oz ol EDG rol Lol via tu effect. Br. ENG. NO ENG EDG CH₂ EDG via via +I +I effec

Since, all the molecules contain same EWG, we can neglect that for now.

Also, Mesomeric effect is stronger than inductive effect.

So, Molecule D will be most reactive and Molecule B (containing both EWG) will be least reactive.

Now, between molecule A and C, C will be more reactive because it contains 2 EDG and also Br is not directly attached to the ring, so its effect will be less.

Therefore, order of reactivity is D > C > A > B.

4.

4. The given Reaction follows Benzyne Mechanism. HE b C Br. NH₂ LCH₂ concerted Mechanism Нco нсон NH3 at (b) NH3 at (@) NH3 N